# Counterfactual Tuning for Temporal Sensitivity Enhancement in Large Language Model-based Recommendation

### v1

@@ 5.1.4. Implementation Details
For traditional methods, we use one self-attention blocks in SASRec and one GRU layer in GRU4Rec, other settings follow the original papers. For LLM-based models, we select Llama3-8B (Dubey et al., 2024) as the backbone and use the AdamW (Loshchilov and Hutter, 2019) optimizer with a learning rate of $1\times 10^{-4}$. Specifically, we implement Llama-based P5 in OpenP5 paper (Xu et al., 2024; Hua et al., 2023). For CETRec, we adopt the two position embedding schemes introduced earlier—SinPE and RoPE—as our implementations of temporal embeddings. Besides, we tune the $\lambda$, the hyperparameter controlling the weight of counterfactual loss in our CT task, in the range $\{0.01,0.1,0.5,1,10\}$. In inference stage, we follow the process in CFT paper (Zhang et al., 2024b). We generate five items, finding the most closely matched actual item for each of them to generate the Top-5 recommendation list. The Top-10 recommendation list is generated from the second-best matched items of each item. All experiments are conducted with Python 3.12, PyTorch 2.2.2 and transformers 4.43.4.
@@
MovieLens Steam
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
SinPE CETRec 0.1101 0.0693 0.1209 0.0728 0.0670 0.0436 0.0779 0.0471
w.o. CT 0.1035 0.0634 0.1167 0.0676 0.0620 0.0407 0.0720 0.0440
RoPE CETRec 0.1283 0.0779 0.1365 0.0806 0.0888 0.0635 0.0997 0.0670
w.o. CT 0.1192 0.0745 0.1300 0.0781 0.0737 0.0474 0.0821 0.0501
w.o. Both 0.0985 0.0614 0.1093 0.0650 0.0595 0.0408 0.0687 0.0438
@@ 5.2. Performance Comparison(RQ1 & RQ2)
  • CETRec with RoPE outperforms all the baseline models on two datasets except HR@10 on Steam, achiving an average relative improvement of 11.6% across all metrics and datasets. This shows the effectiveness of CETRec.
@@ 5.4. Temporal Sensitivity Analysis(RQ4)
We further conduct experiments to investigate the temporal sensitivity of CETRec. By reversing the interaction sequences, we alter both the absolute order and the relative order of each sequence, thereby changing the temporal information they contain. Therefore, we evaluate the model’s performance on the reversed interaction sequences in the test set and quantify its temporal sensitivity by measuring the change rate to its performance on the original test set. We use the MovieLens dataset as an example. The results are shown in Figure 4, and we draw the following conclusions:
@@ 6. Conclusion
In this paper, we identify and analyze a critical limitation of LLM-based sequential recommendation systems: their inability to fully leverage the rich temporal information embedded in user interaction sequences. This limitation stems from fundamental architectural constraints in Transformer-based models, where self-attention mechanisms lack inherent sequence ordering and position embeddings designed for natural language prove inadequate for modeling user preference evolution. To address this challenge, we propose Counterfactual Enhanced Temporal Framework for LLM‑Based Recommendation (CETRec), a novel approach grounded in causal inference principles. CETRec first conceptualizes temporal order as an independent causal factor in recommendation outcomes, distinct from item content. By applying temporal embeddings at the item level and constructing a counterfactual world where temporal information is erased, we design a specialized tuning objective that directly optimizes the model’s temporal sensitivity. Our experiments across two real-world datasets demonstrate that CETRec consistently outperforms state-of-the-art baselines. Ablation studies confirm that both temporal embedding and counterfactual tuning components contribute significantly to performance gains, while our temporal sensitivity analysis validates CETRec’s enhanced ability to capture both absolute and relative ordering information.
Looking forward, this work opens several promising research directions. While our current implementation assigns temporal embeddings based solely on interaction order, future work could explore more sophisticated strategies—such as incorporating actual time intervals between interactions, modeling seasonal patterns in user behavior, or integrating collaborative filtering signals—to better capture the complex temporal dynamics in recommendation scenarios. Additionally, extending our counterfactual inference framework to other aspects of recommendation systems, such as user context or item popularity, could yield further improvements in recommendation quality and explainability.
7. GenAI Usage Disclosure
LLM is used to polish the writing of the paper.

### v2

@@ 5.1.4. Implementation Details
For traditional methods, we use one self-attention blocks in SASRec and one GRU layer in GRU4Rec, other settings follow the original papers. For LLM-based models, we select Llama3-8B (Dubey et al., 2024) as the backbone and use the AdamW (Loshchilov and Hutter, 2019) optimizer with a learning rate of $1\times 10^{-4}$. Specifically, we implement Llama-based P5 in OpenP5 paper (Xu et al., 2024; Hua et al., 2023). For CETRec, we adopt the two position encoding schemes introduced earlier—SinPE and RoPE—as our implementations of temporal embeddings. Besides, we tune the $\lambda$, the hyperparameter controlling the weight of counterfactual loss in our CT task, in the range $\{0.01,0.1,0.5,1,10\}$. In inference stage, we follow the process in CFT paper (Zhang et al., 2024b). We generate five items, finding the most closely matched actual item for each of them to generate the Top-5 recommendation list. The Top-10 recommendation list is generated from the second-best matched items of each item. All experiments are conducted with Python 3.12, PyTorch 2.2.2 and transformers 4.43.4.
@@
MovieLens Steam LastFM
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
SinPE CETRec 0.1101 0.0693 0.1209 0.0728 0.0670 0.0436 0.0779 0.0471 0.0780 0.0520 0.0910 0.0550
w.o. CT 0.1035 0.0634 0.1167 0.0676 0.0620 0.0407 0.0720 0.0440 0.0721 0.0482 0.0865 0.0514
RoPE CETRec 0.1283 0.0779 0.1365 0.0806 0.0921 0.0641 0.1072 0.0690 0.0866 0.0591 0.1030 0.0646
w.o. CT 0.1192 0.0745 0.1300 0.0781 0.0856 0.0613 0.1021 0.0669 0.0805 0.0565 0.0980 0.0626
w.o. Both 0.0985 0.0614 0.1093 0.0650 0.0778 0.0530 0.0979 0.0596 0.0264 0.0192 0.0356 0.0221
@@ 5.2. Performance Comparison(RQ1 & RQ2)
  • CETRec with RoPE outperforms all the baseline models on three datasets, achiving improvement across all metrics and datasets. This shows the effectiveness of CETRec.
@@ 5.4. Temporal Sensitivity Analysis(RQ4)
We further conduct experiments to investigate the temporal sensitivity of CETRec. By reversing the interaction sequences, we alter both the absolute order and the relative order of each sequence, thereby changing the temporal information they contain. Therefore, we evaluate the model’s performance on the reversed interaction sequences in the test set and quantify its temporal sensitivity by measuring the change rate to its performance on the original test set. BIGRec and CFT are considered as baselines because of their similar paradigm, effectively demonstrating the impact of different methods on temporal sensitivity of models. We use the MovieLens dataset as an example. The results are shown in Figure 4, and we draw the following conclusions:
@@
5.5. Case Sduty
Figure 5. Case study on Steam Dataset.
[…]
Case study on Steam Dataset.
To illustrate the effectiveness of CETRec and its ability to exploit temporal information in interaction sequences, we conduct a case study on the Steam dataset, comparing SASRec, BIGRec, and CETRec with RoPE. As shown in Figure 5, for a user who sequentially played ”The Next Door”, ”Press X to Not Die”, ”What’s Under Your Blanket!?”, ”2064: Read Only Memories”, ”Incoming Forces”, ”ABZU”, ”Corrosion: Cold Winter Waiting [Enhanced Edition]”, ”Kraven Manor”, ”Morphine”, and ”Transmissions: Element 120”: SASRec recommended ”Layers of Fear”, following the genres of psychological horror adventure games. BIGRec considered the content of the above games and recommended ”The Dweller”. In contrast, CETRec captured the user’s evolving interest in both horror adventure and puzzle games over time, and correctly recommended ”A Girls Fabric Face”.
@@ 6. Conclusion
In this work, we identify the limited temporal awareness of LLM-based sequential recommenders and propose CETRec, a causal inference–driven framework that treats temporal order as an independent causal factor. By applying item-level temporal embeddings and counterfactual tuning, CETRec significantly improves temporal sensitivity and achieves consistent gains over mutipule baselines on real-world datasets. Future work could extend this framework by incorporating richer temporal signals (e.g., time intervals, seasonal patterns) and applying counterfactual reasoning to other recommendation factors such as user context or item popularity.